OpenStudy (anonymous):

How can you say that the function is maximum or minimum? i have a solution then check if its correct....

OpenStudy (anonymous):

y=2+12x-x^3 =12-3x^2 =3(x^2-4) = (x+2)(x-2) x+2=0 x-2=0 x=-2 x=2 a=2 x,2;x=1 y=(x+2)(x-2) =(3)(-1) y′ =-3 x.2;x=3 y=(3=2)(3-2) y′ =5 x=a;x=2 y=(2+2)(2-2) y′ =0 so x=2 at minimum?

jimthompson5910 (jim_thompson5910):

min is at x= -2 max is at x = 2

OpenStudy (anonymous):

@ jim..can you explain how did you come up with that?pls

jimthompson5910 (jim_thompson5910):

y=2+12x-x^3 y'=-3x^2+12 y''=-6x Set first derivative equal to zero and solve to get 0 = -3x^2+12 -12 = -3x^2 4 = x^2 x^2 = 4 x = 2 or x = -2 So we know that at those x values, the tangent line is horizontal, therefore, those points are either minimums or maximums To figure out which is which, plug in each extrema into the second derivative y'' = -6x so when x = 2, y'' = -6(2) = -12 Since this is less than zero, this means that we have a max here when x = -2, y'' = -6(-2) = 12 because this is greater than zero, we have a min here

OpenStudy (anonymous):

thanks jim

jimthompson5910 (jim_thompson5910):

which part doesn't make sense?

OpenStudy (anonymous):

your using 2nd derivative test right?while i'm using 1st derivative test.but the answer will be the same?

jimthompson5910 (jim_thompson5910):

yes either method will get you the same answer, but the first derivative test takes a bit more work

jimthompson5910 (jim_thompson5910):

with the first derivative test, you have to examine the change in sign of the derivatives

OpenStudy (anonymous):

ok tnx.i'll try to use the 2nd...

jimthompson5910 (jim_thompson5910):

whereas with the second derivative test, all you need is the sign itself

jimthompson5910 (jim_thompson5910):

things to remember if f''(x) < 0, then the function is concave down at that point if f''(x) > 0, then the function is concave up at that point

OpenStudy (anonymous):

@ jim..can you explain how did you come up with that?pls

OpenStudy (anonymous):

tnx again...